OpenStudy (anonymous):

Use elimination to solve the system. What is the value of x? 4x + 7y = -14 and 8x + 5y = 8 Answer 3.5 -1.5 8 -4

OpenStudy (anonymous):

multiplying eqn by 2, 8x+14y=-28 now subtracting, 8x+14y=-28 8x+5y=8 9y=-36 y=-4

OpenStudy (anonymous):

4x+7*-4=-14 4x-28=-14 4x=-14+28 4x=14 x=14/4

OpenStudy (anonymous):

3.5